OpenStudy (anonymous):

In an experiment, a petri dish with a colony of bacteria is exposed to cold temperatures and then warmed again. a. Find a quadratic model for the data in the table. b. Use the model to estimate the population of bacteria at 9 hours. Time (hours) 0 1 2 3 4 5 6 Population (1000s) 5.1 3.03 1.72 1.17 1.38 2.35 4.08

OpenStudy (anonymous):

@JakeV8

OpenStudy (anonymous):

Do you have a graphing calculator ?

OpenStudy (anonymous):

yes but i need the work

OpenStudy (anonymous):

The general formula for quadratic eqn is y = ax^2 + bx + c

OpenStudy (anonymous):

take any 3 points from the table (0, 5.1) (1, 3.03) (5, 2.35) plug these points into the above equation and you will get 3 different equations use substitution or any method that you know to solve these 3 eqns with 3 unknowns

OpenStudy (anonymous):

i dont know how to plug points in

OpenStudy (anonymous):

ive been on this test for 6 hours i have 15 min left to answer three questions

OpenStudy (anonymous):

a point is written in the form (x,y) so plug 0 for x and 5.1 for y to get 1st eqn. follow same method to get other 2 eqns

OpenStudy (anonymous):

what

OpenStudy (anonymous):

i have 14 minuetes left to answer my last three questions can someone hurry up and explain this

OpenStudy (anonymous):

5.1 = a*0^2 + b*0 + c 3.03 = a*1^2 + b*1 + c 2.35 = a*5^2 + b*5 + c

OpenStudy (anonymous):

now what is this

OpenStudy (anonymous):

solve of a, b and c

OpenStudy (anonymous):

omg screw this im skipping this question again and going to my other 2 questions

OpenStudy (anonymous):

1tell me about it i hate ashworth @jennyjewell25 lol

Directrix (directrix):

At this link is what may be the same problem. See what you think. http://staff.bbhcsd.org/bradacm/files/2013/10/Enriched-alg-2-trig-unit-2-test-review-key.pdf